OpenStudy (anonymous):

Hi, I'm new and really struggling with this question: Monochromatic light with a wavelength of 6.4 E -7 meter passes through two narrow slits, producing an interference pattern on a screen 4.0 meters away. The first order bright band lines are 2.0 E -2 meters away from the central bright maxima. What is the distance between the slits?

OpenStudy (michele_laino):

That is the Young interference. So we have to apply this formula: \[\large d\sin \varphi = \lambda \] where d is the requested distance, and: \[\large \sin \varphi \cong \frac{{2 \times {{10}^{ - 2}}}}{4} = ...?\]

OpenStudy (michele_laino):

and \[\large \lambda \] is the wavelength of the employed radiation